OpenStudy (anonymous):

w = -4xy + yz - 2xz, x = st, y = e^{st}, z = t^2 what is partial w/partial t} at (5, -2)

OpenStudy (anonymous):

s=5 and t=-2

OpenStudy (anonymous):

-20e^-10 -40 + 220e^-10

OpenStudy (anonymous):

is what i got but its wrong

OpenStudy (anonymous):

\[Find \frac{ dx }{ dt },\frac{ dy }{dt },\frac{ dz }{dt } ,and substituting \frac{ dw }{dt }\] and finally put s=5,x=5,y=-2

OpenStudy (anonymous):

how did u get the x and y and z

OpenStudy (anonymous):

\[\frac{ dx }{ dt }=s,\frac{ dy }{ dt }=e ^{st}*s,\frac{ dz }{ dt }=2t,find \frac{ dw }{dt }\]

OpenStudy (anonymous):

\[\frac{ d }{dt }\left( x y \right)=x \frac{ dy }{dt }+y \frac{ dx }{dt }\]

OpenStudy (loser66):